OpenStudy (driftracer305):

A sporting good store received an order of 80 baseball caps. of which 12 were green. If 1 of the 80 caps is selected at random, what is the probability it will not be green?

OpenStudy (driftracer305):

@ganeshie8 ?

OpenStudy (driftracer305):

@jim_thompson5910 ..... uhh help.....please

OpenStudy (ikram002p):

well whats the probability that its green ?

OpenStudy (driftracer305):

i guess 62/80.....

OpenStudy (driftracer305):

sorry 12/80*

OpenStudy (driftracer305):

uhh ......... i meant not green are 62/80..... but wat to do to get it in percent form?

OpenStudy (ikram002p):

yes ur correct its 12/80 nw the probability of having not green is 1- the probability of having a green

OpenStudy (driftracer305):

0.88?

OpenStudy (ikram002p):

wait its not 62 ::O how did u get 62 ?

OpenStudy (driftracer305):

question says what is the probability that it will NOT be green?

OpenStudy (ikram002p):

well as i said probability of having green is 12/80 = 0.15 the probability of having NOT green is 1- 0.15 got it nw ?

OpenStudy (driftracer305):

ohh 0.85..... yea got it boss......... thx

OpenStudy (ikram002p):

np :)

OpenStudy (driftracer305):

@aum ... can i ask u a question?

OpenStudy (aum):

yes?

OpenStudy (driftracer305):

so he got 0.15 by 12/80......... but y cant we do 62/80 for the probability that is not green?

OpenStudy (aum):

because 80 - 12 = 68 (not 62)

OpenStudy (driftracer305):

ahh sorry my bad..... i meant 68.... sorry....

OpenStudy (aum):

68/80 = 17/20 = 0.85

OpenStudy (driftracer305):

so we can do that as well right? 68 /80 instead of 12/80 then subtract it from 1?

OpenStudy (aum):

Yes.

OpenStudy (driftracer305):

ohh ok kool....... makes a lot of sense now..........thx a lot

OpenStudy (aum):

you are welcome.

OpenStudy (oleg3321):

please help with question @driftracer305

OpenStudy (driftracer305):

k wat is it?